OpenStudy (anonymous):

Help please ..... How far of a small sphere of 100 kg a particle would have to be placed so that the ball would attract the particle with the same force as the Earth? This experiment could be done effectively? How come?

OpenStudy (help!!!!):

How far from the earth?

OpenStudy (anonymous):

the distance the question does not specify

OpenStudy (abhisar):

Mass of particle ?

OpenStudy (fifciol):

 \[\frac{ Gm_p m_s}{r_1^2 }=\frac{ GM_e m_s}{r_2^2 } \Rightarrow r_2 =r_1\sqrt{\frac{ M_e }{m_s }}\]

OpenStudy (fifciol):

the distance from particle to the Earth is sqrt Me/ms larger than the distance from particle to sphere, with these givens I'm not sure there could be more accurate solution

OpenStudy (mrnood):

I think you have misinterpreted the question. My interpretation is Take the weight of the particle on earth What distance would a 100kg mass have to be to exert the same force (in the ABSENCE of other forces) Weight = GMeM1/re^2 (re = radius of earth) Force = G 100 M1/d^2 (d = distance from mass to particle) d^2/re^2 = G100M1/GMeM1 =100/Me SO d^2 = 100 re^2/Me d= re * sqrt (100/Me)

OpenStudy (mrnood):

A quick estimate - I make this approx. 1x10-4m = 0.1mm You could not get a point mass of 100kg so small....

OpenStudy (anonymous):

I agree with MrNood

OpenStudy (anonymous):

correct????

OpenStudy (mrnood):

Are you saying this is the solution? It looks about right to me, and corresponds with the statement that oyu cannot get a point mass of sufficient denity.